OpenStudy (anonymous):

solve In 2+ In x= 5 I don't understand please help WILL GIVE MEDALS

OpenStudy (mathmale):

There are 3 important rules for logs in problems such as this one. ln a + ln b = ln (ab) ln a - ln b + ln (a/b) b*ln a = ln (a^b) May I assume you're familiar with these? If so, which one applies to the problem you've posted?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

OK. Right. If ln a + ln b = ln (ab), then ln 2 + ln 5 = ln (???)

OpenStudy (mathmale):

How are you doing with this problem so far, DMB?

OpenStudy (mathmale):

DMB: I surely would have liked to have helped you to find the solution to this problem. Next time, please let me know before you log off.

OpenStudy (anonymous):

ln 2 + ln x = 5. Per the rules of log we convert to ln (2x) = 5. Now you must expointiate the log. Since it is a natural log we use e. e^ln(2x) = e^5. This will give you 2x = e^5. Solve from there.

OpenStudy (loser66):

@mathmale I think he was kicked out of the net so that he logged off without warning you.

OpenStudy (anonymous):

Hey @mathmale sorry got kick off and then had to go to tennis practice

OpenStudy (anonymous):

@nelsonjedi how do i solve that

OpenStudy (amoodarya):



OpenStudy (anonymous):

is that the answer?

OpenStudy (amoodarya):

yes but it is 2x so divide by 2

OpenStudy (anonymous):

how do you do that

OpenStudy (amoodarya):

\[\log_{b} A +\log_{b} D =\log_{b} AD \]

OpenStudy (amoodarya):

\[\ln x =\log_{e} x\] \[\log_{e} x+\log_{e} 2 =5\rightarrow \log_{e} 2x =5 \rightarrow \log_{e} 2x=5\log_{e} e\]

OpenStudy (amoodarya):

\[\log_{e} 2x=\log_{e} e ^{5} \rightarrow 2x=e ^{5} \rightarrow x=\frac{ 1 }{ 2 }e ^{5}\]

OpenStudy (anonymous):

so the answer is \[\frac{ 1 }{ 2 }e ^{5}\]

OpenStudy (amoodarya):

yes

OpenStudy (amoodarya):

understand now ?

OpenStudy (anonymous):

yes thank you